OpenStudy (lgbasallote):

Given: \[A = \{x: 1 \le x \le 5\}\] why is \[A' = \{x : x < 1 \vee x > 5\}\] and not \[A' = \{x: x< 1 \wedge x > 5\}\]

OpenStudy (anonymous):

Since \(x\) cannot be both less than one and greater than five at one time (otherwise we have a logical contradiction). Also, De Morgan's laws are useful for this.

OpenStudy (lgbasallote):

so does that mean \[A = \{x : x \ge 1 \wedge x \le 5\}\]

OpenStudy (anonymous):

In any case: \[ A=\{x|1\leq x\leq 5\}=\{x|1\leq x \wedge x\leq 5\} \]

OpenStudy (anonymous):

Yessir.

OpenStudy (lgbasallote):

i can't believe i didn't think of that first reason....

OpenStudy (anonymous):

Unless your x's can quantum tunnel . . .

OpenStudy (lgbasallote):

uh-oh....math conversation.....time to leave..

OpenStudy (anonymous):

(Sorry, I've been hanging out in the Physics section too long)

OpenStudy (anonymous):

Go go, quantum entanglement and random transmission of data!

OpenStudy (anonymous):

Strange things can happen in this here multiverse.